4.3.54 Question Help The chance of winning a lottery game is 1 in approximately 21 million. Suppose you buy a $1 lottery tickNeed second part!! thanks
Answer #1

We are calculating the expected value

= xP(X)

We can win 9,000,000 with probability 1 in 21m therefore p = 1/21m

Therefore the chance of losing = 1 - 1/21m

But we certainly spend $1. therefore even if we win the net winning amt = 9m -1

Losing amount = -$1

The expected win = 1 /21m *( 9m -1)+ (1 - 1/21m) * (-1)

= -$0.57

This means that on an avergae our net winning will be -$0.57 per game.

First blank = Net winning (it can be loss then the 2nd blank can be 0.57

Second blank = -0.57

(it can be loss then the 2nd blank can be 0.57)

(it can be profit then the 2nd blank can be -0.57)